Q4) Code Analysis (10 Points) What do the following loops print? Work out the answer by tracing the code, not by using the computer a. (2 Points) b. (2 Points) c. (2 Points) for n in range (1, 6): for n in range(1, 11): for n in range(1, 6): n=n+2 print(s) print(s) print (s, n) Rewrite the following for loop as a while loop (4 Points). fori in range(1, 10): Instructions Use only the concepts from Chapter 2 thru 5 Python
Answer #1

Hi, you will need to understand the meaning of range function to answer this,

range(start ;stop; step) which says interate from start till stop(excluding stop) incrementing each time by step

which is roughly equivalent to for(i=start;i<stop;i=i+step)

1. so in 1st question, no step is given, default is 1.

so the loop prints 5 numbers (1-6) for n=1,2,3,4,5

n=1: s=s+n=1+1= 2

n=2: s=s+n=2+2= 4

n=3: s=s+n=4+3= 7

n=4: s=s+n=7+4= 11

n=5: s=s+n=11+5= 16

2.this is almost same as 1st question, but the interesting part is loop variable is updated inside the loop,

but in pyhton, for loop iterates over all the numbers in range() so even if n was updated as soon as we go for next iteration it is assigned the next value from range(1,11). hence it prints,

4
8
13
19
26
34
43
53
64
76

3. this one follows same concept as 2 and infact is exactly same as question 1 output wise because n has the same range, only difference being print is outside of for loop, so only printed once,

hence prints (16,6)

Now, to tackle the above situations we can use while loops,

we are asked to rewrite

s=0

for i in range(1,10):

s=s+i

which essentially adds in incremntal value to s and finally at 9 we get s= 46

with while, we explicitly have to declare i and increment it like this,

s=1
i=1
while i<10:
    s=s+i
    i+=1
    print(s)

Thumbs up if this was helpful, otherwise let me know in comments. Good day.